a car covers the first 39km of its journey in 45 minutes and the remaining 25km in 35 minutes. what is the average speed of the car???? please give me the solution

Answer by edjones(8007)

(39+25)/(45+35)=64/80
x/60=64/80
x=3840/80
x=48 k/h Speed of the car.
.
Ed